3. a step-up transformer increases 23 V to 120 V. what is the current in the...

3.
a step-up transformer increases 23 V to 120 V.
what is the current in the secondary coil as compared to the primary coil? assume 100% efficiency.
Answer #1

Let Vp be the Voltage through the primary coil

Ip be the current through the primary coil

Vs be the voltage through the secondary coil

Is be the current through the secondary coil.

Now, 100 percent efficiency

So Power of primary = Power of secondary coil

Hence, Vp * Ip = Vs* Is

So, Is/Ip = Vp/Vs = 23/120 = 0.1917

Hence Is = 0.1917* Ip